Find the exponential decay function for a radioactive material that has a half-life of 1234 years. How long will it take until only 10% of the material remains.
P(T) = P e^-kt

Answer by stanbon(48535)

Find the exponential decay function for a radioactive material that has a half-life of 1234 years. How long will it take until only 10% of the material remains.
----------------
P(T) = P e^-kt
0.10P = P*e^(-t/1234)
0.10 = e^(-t/1234)
Take the natural log of both sides to get:
-t/1234 = ln(0.10)
-t = 1234*-2.3026
t = 2841.39 years
---------------------
Cheers,
Stan H.